Circular No. 8706

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 H1 (McNAUGHT) R. H. McNaught reports his discovery on a comet on CCD images taken

with the 0.5-m Uppsala Schmidt (discovery observation tabulated below) in the course of the Siding Spring Survey, the object having an asteroidal head with a very weak, diffuse tail 8′′ long in p.a. ∼ 250^{o}. Following posting on the ‘NEO Confirmation Page’, A. C. Gilmore reports that his CCD images taken on Apr. 30.7 UT with the 1.0-m f/7.7 reflector at Mount John through cirrus clouds show a small, condensed head with a narrow tail to the west-southwest. Images taken by J. Young with the Table Mountain 0.6-m Cassegrain reflector on May 1.5 in twilight show a round 6′′ coma with no noticeable central condensation and no tail.

2006 UT  α_{2000}  δ_{2000}  Mag.
Apr. 29.78705  22^{h}29^{m}25.^{s}44  −19^{o}35′46′′.1  17.9

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2006-J06.

T = 2006 Feb. 24.566 TT  ω = 285.^{o}416
e = 0.38667  Ω = 357.078 }2000.0
q = 2.15508 AU  i = 13.045
a = 3.51371 AU n^{o} = 0.149642 P = 6.6 years

V5117 SAGITTARII D. K. Lynch, R. W. Russell, and R. J. Rudy, The Aerospace Corpora-

tion; and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of V5117 Sgr (cf. IAUC 8673) on May 1 UT using SpeX at the Infrared Telescope Facility. The nova is a classical Fe II nova that is still in the low-excitation stage. It showed strong emission of the Brackett and Pashen lines, the Lyβ-fluoresced O I lines, C I and N I, Fe II emission, and weak He I 1.0830- and 2.0581-μm lines. There was also a strong continuum increasing to longer wavelengths, indicative of thermal emission from dust. The thermal continuum was well fitted by an 1100-K black body. E(B −V) = 0.50 ± 0.15 was obtained from the O I lines. Infrared magnitudes (± 0.1): J = 10.6, H = 9.5, K = 7.9.

Visual magnitude estimates by A. Pearce, Nedlands, W. Australia: Feb. 20.842 UT, 10.1; 21.844, 10.1. CCD red-magnitude from D. Mendicini, Santa Fe, Argentina (provided by E. Waagen, AAVSO): Mar. 24.208, 10.5.

2006 May 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green